# ZMP-HH-25/5

### v1

@@ Abstract
[…]
@@ 1 Introduction
However, establishing the validity of the Emergent String Conjecture in this setting requires information beyond the one accessible from a purely algebraic perspective. First, it is crucial to establish the criticality of the emergent string, as well as its uniqueness. Furthermore, demonstrating the existence of a tower of BPS particles requires a careful analysis of the BPS indices associated with special Lagrangian 3-cycles, which is a notoriously complicated problem [53]. Finally, to properly unveil the right duality frame that emerges in the infinite-distance limit, it is paramount to understand the worldsheet theory of the emergent string. To tackle these challenges, it is clear that a proper geometric characterization of the underlying Calabi–Yau manifold in type II limits is of utmost importance. This is similar in spirit to the geometric analysis of complex structure limits for Weierstrass models [12, 13, 15, 16] in the context of F-theory compactifications. To this end, we will focus our attention on a particular class of type II limits which are realised by a so-called Tyurin degeneration [54], of which the aforementioned K-points [33] are a special case. [^2] In Section 6 we address type II limits which do not correspond to Tyurin degenerations. For an illustration, see Figure -957. These are limits in the complex structure moduli space in which the Calabi–Yau threefold splits into two components that intersect over a K3 surface, such that complete control over the underlying geometry is ensured. In the context of mirror symmetry, these types of degenerations have been studied in the physics literature in [55, 56], but we stress that they can also occur away from the large complex structure point, as exemplified by the K-point.
@@
{mdframed}
[backgroundcolor=white,shadow=true,shadowsize=4pt,shadowcolor=black,roundcorner=6pt] In Type IIB string theory compactified on Calabi–Yau threefolds, a unique emergent, critical heterotic string with perturbative gauge group of rank $2+b$ becomes tensionless at any Tyurin degeneration realising a type IIb limit in the complex structure moduli space of the threefold. The string is always accompanied by a tower of BPS states corresponding to winding and momentum modes in accordance with the Emergent String Conjecture.
@@
We also compare our findings with previous results concerning emergent strings arising in the Kähler moduli space of Type IIA string theory on K3-fibred Calabi–Yau threefolds [8]. In this setting, the emergent heterotic string arises from NS5-branes wrapping the generic K3-fibre $\widehat{Z}$ of the threefold, and taking the large-volume limit of the base while keeping the volume of the generic fibre finite. By comparing the worldsheet theories at various levels of detail, we find that they have the same massless spectrum, and that the number of free fields is the same if the two underlying K3s $Z$ and $\widehat{Z}$ are mirror to each other. Additionally, we argue that in setups with a perturbative heterotic dual string, the interactions on the two worldsheet theories agree if the underlying Calabi–Yau threefolds are mirror to each other. Most notably, in a particular setting where we can identify a special Lagrangian $T^{3}$-fibration such that one can apply the SYZ conjecture, we explicitly show that the geometries on the IIA and IIB side are mapped to each other under mirror symmetry,
@@ 2 Type II Degenerations of Calabi–Yau Threefolds
The situation is depicted in Figure -959 for $r=2$.
[…]
Figure -959: A local patch in a complex two-dimensional vector multiplet moduli space $\mathcal{M}_{\mathrm{V}}$ containing two divisors $\Delta_{1}$ and $\Delta_{2}$ which intersect at a point $\Delta_{12}$.
@@ 2.1.1 Limiting Mixed Hodge Structures
It is convenient to package the information contained in the mixed Hodge structure (2.9) into a so-called Hodge–Deligne diamond. Similar to a Hodge diamond, this is a diagram in which the various $I^{p,q}$ spaces are depicted in a square as in Figure -958, with a dot signifying that the corresponding space is non-trivial, and their labels indicating the dimension $i^{p,q}$. More information on the relations between the dimensions $i^{p,q}$ and the resulting Hodge–Deligne diamonds is reviewed in Appendix A.
$b$$b$$b^{\prime}$$b^{\prime}$
Figure -958: The Hodge–Deligne diamond corresponding to a type $\mathrm{II}_{b}$ singularity. The horizontal rows indicated in black, blue, and red correspond to the graded spaces $\mathrm{Gr}_{4}$, $\mathrm{Gr}_{3}$, and $\mathrm{Gr}_{2}$, respectively. Here we put $b^{\prime}=h^{2,1}-b-1$.
@@
[…]
@@
[…]
@@ 2.1.2 Type II enhancement chains and the growth theorem
where we drop the dependence on $\Delta_{k_{1}\cdots k_{r}}$ for notational simplicity, recall also Figure -958. Of special importance will be the space
@@
$\mathrm{II}_{b}\to\mathrm{II}_{b^{\prime}}\,,\qquad\,\,\,\text{for $b\leq b^{\prime}$}$ (2.22)
@@ 2.2 Geometry of Tyurin Degenerations
where we denote the intersection locus of $X_{1}$ and $X_{2}$ by $Z$. For a Tyurin degeneration the manifold $Z$ is a K3 surface. The situation is depicted schematically in Figure -957. To compare the current situation with the setup described in Section 2.1, note that one can equivalently view the fibration (2.27) as an embedding of the disk $\mathbf{D}$ into the complex structure moduli space $\mathcal{M}_{\mathrm{V}}$ of the family $V_{z}$ of Calabi–Yau threefolds. In particular, one may identify the limit $z\to 0$ with some divisor $\{u^{k_{1}}=\dots=u^{k_{r}}=0\}$. To simplify the notation, we will denote the corresponding divisor by $\Delta$ without an additional subscript. [^10] To be clear, this is not to be confused with the notation (2.6) for the full discriminant locus. Before we explain in general why Tyurin degenerations correspond to type II limits in the complex structure moduli space, we provide an illustrative example of a Tyurin degeneration.
@@
Consider the mirror of the Calabi–Yau threefold $\widehat{V}$ obtained as a degree-12 hypersurface in the weighted projected space $\mathbb{P}_{11226}$, denoted by $\mathbb{P}_{11226}[12]$. The Calabi–Yau threefold $\mathbb{P}_{11226}[12]$ with Hodge-numbers $h^{1,1}(\widehat{V})=2,h^{2,1}(\widehat{V})=128$ admits a K3-fibration and is the main example in [8] illustrating the Emergent String Conjecture in the vector multiplet moduli space of Type IIA/M-theory compactifications on Calabi–Yau threefolds. The mirror manifold $V$ to $\widehat{V}$ can be obtained by a Green–Plesser [63] construction. Consider the polynomial
@@
$P\stackrel{{\scriptstyle u_{2}\to 0}}{{\longrightarrow}}u_{2}^{-1}x_{1}^{6}x_{2}^{6}\,,$ (2.30)
@@
Figure -957: A schematic depiction of a Tyurin degeneration (2.27). The fibre $V_{z}$ corresponding to the Calabi-Yau threefold over a generic point in the moduli space is depicted in red, while the central fibre $V_{0}$ is depicted in orange. For a Tyurin degeneration, the latter splits into two components $X_{1}$ and $X_{2}$ which intersect over a K3 surface $Z$, here indicated in blue.
@@
can be thought of as adding a leg normal to $Z$ in $V_{0}$ to a 2-form on $Z$. The first isomorphism states that the elements in $\mathrm{Gr}_{2}(\Delta)$ correspond to 3-forms on the degenerate space $V_{0}$ which can be built from 2-forms on $Z$ by completing them in the direction normal to $Z$ via the map $d^{\ast}$. The second isomorphism identifies this space with the cohomology group of 2-forms on $Z$ which do not arise as the pullback of 2-forms from $X_{1}$ or $X_{2}$ to $Z$. The cohomology of such 2-forms on $Z$ gives rise to the transcendental lattice of the K3 surface $Z$, [^12] For a stand-alone K3 surface, the transcendental lattice is defined as the orthogonal complement of the Picard lattice. The geometry of the 2-cycles associated with the latter is governed by Kähler deformations, whereas the geometry of the 2-cycles associated with the transcendental lattice is governed by complex structure deformations. As the complex structure deformations of a CY threefold are associated with the geometry of 3-cycles, it must be such that the 2-cycles of $Z$ associated with the transcendental lattice must lift to 3-cycles of $V_{0}$.
@@
By (2.18), its signature is [^13] Strictly speaking, at the level of 3-forms the space $\mathrm{Gr}_{2}$ does not immediately come equipped with a polarization, due to the fact that it is not primitive. The space $\mathrm{Gr}_{4}$, however, does carry a polarization which is induced from the polarization (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $V$ together with the action of the log-monodromy operator, see for example [39] for further details. For our purposes, it is sufficient to use the fact that $\mathrm{Gr}_{2}\cong\mathrm{Gr}_{4}$ as vector spaces to speak of the polarization lattice on the K3 surface $Z$.
${\rm sgn}(\Lambda_{\rm trans})=(2,b)\quad\text{in a ${\rm II}_{b}$ Tyurin degeneration.}$. (2.41)
@@ 3.1 BPS-Particle Tower from D3-Branes
$C_{0}=\partial_{*}\iota_{*}\Gamma_{0}\in H_{2}(Z)\,.$ (3.9)
On $\Delta$ we can view $\iota_{*}\Gamma_{0}$ as an $S^{1}$ fibration over $C_{0}$ with the fibre corresponding to the circle that degenerates over $Z$. For the cycle $\Gamma_{0}\in H_{3}(V)$ to be special Lagrangian, $C_{0}$ must lie in the transcendental lattice of $Z$, i.e. the restriction of the Kähler form of $Z$ must vanish, $J_{Z}|_{C_{0}}=0$, and $C_{0}$ is calibrated by the holomorphic 2-form $\Omega^{2,0}$ on $Z$. At least in the degenerate limit of vanishing $S^{1}$ fibre, which we interested in, this is sufficient to guarantee that $\Gamma_{0}$ is special Lagrangian.
[…]
@@ 3.2 Tensionless Strings
where $\partial=(\partial/\partial z){\rm d}z$. Simple BPS solutions to these equations of motions are obtained for holomorphic profiles ¯${\bar{\partial}}t=0$. The Einstein equations then fix the warp factor to be
@@
$\mathcal{I}_{\mathrm{II}}:=\{k\in\{1,\ldots,h^{2,1}\}:\text{$I^{p,q}(\Delta_{k})$ is of type II}\}$ (3.25)
@@
$\frac{T_{\mathbf{e}_{k_{1}}}}{m_{\rm tower}^{2}}\to\text{const.}\,,$. , (3.30)

### v2

@@ Abstract
We study infinite distance limits in the complex structure moduli space of Type IIB compactifications on Calabi–Yau threefolds, in light of the Emergent String Conjecture. We focus on the so-called type II limits, which, based on the asymptotic behaviour of the physical couplings in the low-energy effective theory, are candidates for emergent string limits. However, due to the absence of Type IIB branes of suitable dimensionality, the emergence of a unique critical string accompanied by a tower of Kaluza–Klein states has so far remained elusive. By considering a broad class of type IIb limits, corresponding to so-called Tyurin degenerations, and studying the asymptotic behaviour of four-dimensional EFT strings in this geometry, we argue that the worldsheet theory of the latter describes a unique critical heterotic string on $T^{2}\times\mathrm{K3}$ with a gauge bundle whose rank depends on $b$. In addition, we establish the presence of an infinite tower of BPS particles arising from wrapped D3-branes by identifying a suitable set of special Lagrangian 3-cycles in the geometry. The associated BPS invariants are conjectured to be counted by generalisations of modular forms. As a consistency check, we further show that in special cases mirror symmetry identifies the EFT strings with the well-understood emergent string limits in the Kähler moduli space of Type IIA compactifications on K3-fibred Calabi–Yau threefolds. Finally, we discuss the implications of the Emergent String Conjecture for type II limits which do not correspond to Tyurin degenerations, and predict new constraints on the possible geometries of type II degenerations which resemble those arising in the Kulikov classification of degenerations of K3 surfaces.
@@ 1 Introduction
However, establishing the validity of the Emergent String Conjecture in this setting requires information beyond the one accessible from a purely algebraic perspective. First, it is crucial to establish the criticality of the emergent string, as well as its uniqueness. Furthermore, demonstrating the existence of a tower of BPS particles requires a careful analysis of the BPS indices associated with special Lagrangian 3-cycles, which is a notoriously complicated problem [53]. Finally, to properly unveil the right duality frame that emerges in the infinite-distance limit, it is paramount to understand the worldsheet theory of the emergent string. To tackle these challenges, it is clear that a proper geometric characterization of the underlying Calabi–Yau manifold in type II limits is of utmost importance. This is similar in spirit to the geometric analysis of complex structure limits for Weierstrass models [12, 13, 15, 16] in the context of F-theory compactifications. To this end, we will focus our attention on a particular class of type II limits which are realised by a so-called Tyurin degeneration [54], of which the aforementioned K-points [33] are a special case. [^2] In Section 6 we address type II limits which do not correspond to Tyurin degenerations. For an illustration, see Figure 3. These are limits in the complex structure moduli space in which the Calabi–Yau threefold splits into two components that intersect over a K3 surface, such that complete control over the underlying geometry is ensured. In the context of mirror symmetry, these types of degenerations have been studied in the physics literature in [55, 56], but we stress that they can also occur away from the large complex structure point, as exemplified by the K-point.
@@
In Type IIB string theory compactified on Calabi–Yau threefolds, a unique emergent, critical heterotic string with perturbative gauge group of rank $2+b$ becomes tensionless at any Tyurin degeneration realising a type IIb limit in the complex structure moduli space of the threefold. The string is always accompanied by a tower of BPS states corresponding to winding and momentum modes in accordance with the Emergent String Conjecture.
@@
We also compare our findings with previous results concerning emergent strings arising in the Kähler moduli space of Type IIA string theory on K3-fibred Calabi–Yau threefolds [8]. In this setting, the emergent heterotic string arises from NS5-branes wrapping the generic K3-fibre $\widehat{Z}$ of the threefold, and taking the large-volume limit of the base while keeping the volume of the generic fibre finite. By comparing the worldsheet theories at various levels of detail, we find that they have the same massless spectrum, and that the number of free fields is the same if the two underlying K3s $Z$ and $\widehat{Z}$ are mirror to each other. Additionally, we argue that in setups with a perturbative heterotic dual string, the interactions on the two worldsheet theories agree if the underlying Calabi–Yau threefolds are mirror to each other. Most notably, in a particular setting also considered in [57], a special Lagrangian $T^{3}$-fibration suitable for applying the SYZ conjecture can be identified and we explicitly show that the geometries on the IIA and IIB side are mapped to each other under mirror symmetry.
@@ 2 Type II Degenerations of Calabi–Yau Threefolds
The situation is depicted in Figure 1 for $r=2$.
$u^{1}=0$$u^{2}=0$$\Delta_{1}$$\Delta_{2}$$\Delta_{12}$
Figure 1: A local patch in a complex two-dimensional vector multiplet moduli space $\mathcal{M}_{\mathrm{V}}$ containing two divisors $\Delta_{1}$ and $\Delta_{2}$ which intersect at a point $\Delta_{12}$.
@@ 2.1.1 Limiting Mixed Hodge Structures
It is convenient to package the information contained in the mixed Hodge structure (2.9) into a so-called Hodge–Deligne diamond. Similar to a Hodge diamond, this is a diagram in which the various $I^{p,q}$ spaces are depicted in a square as in Figure 2, with a dot signifying that the corresponding space is non-trivial, and their labels indicating the dimension $i^{p,q}$. More information on the relations between the dimensions $i^{p,q}$ and the resulting Hodge–Deligne diamonds is reviewed in Appendix A.
$bb^{\prime}bb^{\prime}$
Figure 2: The Hodge–Deligne diamond corresponding to a type $\mathrm{II}_{b}$ singularity. The horizontal rows indicated in black, blue, and red correspond to the graded spaces $\mathrm{Gr}_{4}$, $\mathrm{Gr}_{3}$, and $\mathrm{Gr}_{2}$, respectively. Here we put $b^{\prime}=h^{2,1}-b-1$.
@@
The reason why the asymptotic split (2.9) of the space of harmonic 3-forms is important for us is because it allows us to read off the asymptotic behaviour of the physical couplings appearing in the low-energy effective theory (2.2). This is because the so-called growth theorem of [49, 50] provides an asymptotic expression for the Hodge norm $\|\cdot\|$ defined in (2.3). To state the precise result, we introduce coordinates
@@
for some constant $\gamma>1$, and where we set $n=h^{2,1}$. In words, the growth sector $R_{k_{1}\cdots k_{n}}$ parametrises the region in the (local patch of the) moduli space which is sequentially closest to $\Delta_{k_{1}}$, then $\Delta_{k_{2}}$, etc. Furthermore, the parameter $\gamma$ controls the precise hierarchy. Correspondingly, the growth theorem provides an asymptotic expression for the Hodge norm $\|\cdot\|$ with corrections being of order $\gamma^{-1}$.
@@ 2.1.2 Type II enhancement chains and the growth theorem
where we drop the dependence on $\Delta_{k_{1}\cdots k_{r}}$ for notational simplicity, recall also Figure 2. Of special importance will be the space
@@
[…]
@@ 2.2 Geometry of Tyurin Degenerations
where we denote the intersection locus of $X_{1}$ and $X_{2}$ by $Z$. For a Tyurin degeneration the manifold $Z$ is a K3 surface. The situation is depicted schematically in Figure 3. To compare the current situation with the setup described in Section 2.1, note that one can equivalently view the fibration (2.27) as an embedding of the disk $\mathbf{D}$ into the complex structure moduli space $\mathcal{M}_{\mathrm{V}}$ of the family $V_{z}$ of Calabi–Yau threefolds. In particular, one may identify the limit $z\to 0$ with some divisor $\{u^{k_{1}}=\dots=u^{k_{r}}=0\}$. To simplify the notation, we will denote the corresponding divisor by $\Delta$ without an additional subscript. [^10] To be clear, this is not to be confused with the notation (2.6) for the full discriminant locus. Before we explain in general why Tyurin degenerations correspond to type II limits in the complex structure moduli space, we provide an illustrative example of a Tyurin degeneration.
@@
Consider the mirror of the Calabi–Yau threefold $\widehat{V}$ obtained as a degree-12 hypersurface in the weighted projected space $\mathbb{P}_{11226}$, denoted by $\mathbb{P}_{11226}[12]$. The Calabi–Yau threefold $\mathbb{P}_{11226}[12]$ with Hodge-numbers $h^{1,1}(\widehat{V})=2,h^{2,1}(\widehat{V})=128$ admits a K3-fibration and is the main example in [8] illustrating the Emergent String Conjecture in the vector multiplet moduli space of Type IIA/M-theory compactifications on Calabi–Yau threefolds. The mirror manifold $V$ to $\widehat{V}$ can be obtained by a Greene–Plesser [64] construction. Consider the polynomial
@@
$P\mathrel{{\mathop{\longrightarrow}\limits^{u_{2}\to 0}}}u_{2}^{-1}x_{1}^{6}x_{2}^{6}\,$, (2.30)
@@
Figure 3: A schematic depiction of a Tyurin degeneration (2.27). The fibre $V_{z}$ corresponding to the Calabi-Yau threefold over a generic point in the moduli space is depicted in red, while the central fibre $V_{0}$ is depicted in orange. For a Tyurin degeneration, the latter splits into two components $X_{1}$ and $X_{2}$ which intersect over a K3 surface $Z$, here indicated in blue.
@@
can be thought of as adding a leg normal to $Z$ in $V_{0}$ to a 2-form on $Z$. The first isomorphism states that the elements in $\mathrm{Gr}_{2}(\Delta)$ correspond to 3-forms on the degenerate space $V_{0}$ which can be built from 2-forms on $Z$ by completing them in the direction normal to $Z$ via the map $d^{\ast}$. The second isomorphism identifies this space with the cohomology group of 2-forms on $Z$ which do not arise as the pullback of 2-forms from $X_{1}$ or $X_{2}$ to $Z$. The cohomology of such 2-forms on $Z$ gives rise to the transcendental lattice of the K3 surface ZZ, [^12] For a stand-alone K3 surface, the transcendental lattice is defined as the orthogonal complement of the Picard lattice. The geometry of the 2-cycles associated with the latter is governed by Kähler deformations, whereas the geometry of the 2-cycles associated with the transcendental lattice is governed by complex structure deformations. As the complex structure deformations of a CY threefold are associated with the geometry of 3-cycles, it must be such that the 2-cycles of $Z$ associated with the transcendental lattice must lift to 3-cycles of $V_{0}$.
@@
By (2.18), its signature is [^13] Strictly speaking, at the level of 3-forms the space $\mathrm{Gr}_{2}$ does not immediately come equipped with a polarization, due to the fact that it is not primitive. The space $\mathrm{Gr}_{4}$, however, does carry a polarization which is induced from the polarization $(\cdot,\cdot)$ on $V$ together with the action of the log-monodromy operator, see for example [39] for further details. For our purposes, it is sufficient to use the fact that $\mathrm{Gr}_{2}\cong\mathrm{Gr}_{4}$ as vector spaces to speak of the polarization lattice on the K3 surface $Z$.
${\rm sgn}(\Lambda_{\rm trans})=(2,b)\quad\text{in a ${\rm II}_{b}$ Tyurin degeneration.}$ (2.41)
@@ 3.1 BPS-Particle Tower from D3-Branes
$C_{0}\in H_{2}(Z)$ (3.9)
with the following property: On $\Delta$ we can view $\iota_{*}\Gamma_{0}$ as an $S^{1}$ fibration over $C_{0}$ with the fibre corresponding to the circle that degenerates over $Z$. For the cycle $\Gamma_{0}\in H_{3}(V)$ to be special Lagrangian, $C_{0}$ must lie in the transcendental lattice of $Z$, i.e. the restriction of the Kähler form of $Z$ must vanish, $J_{Z}|_{C_{0}}=0$, and $C_{0}$ is calibrated by the holomorphic 2-form $\Omega^{2,0}$ on $Z$. At least in the degenerate limit of vanishing $S^{1}$ fibre, which we interested in, this is sufficient to guarantee that $\Gamma_{0}$ is special Lagrangian.
To show the existence of a tower of states, we now consider multi-wrappings of D3-branes on $\Gamma_{0}$. By wrapping $n$ D3-branes on $\Gamma_{0}$ we obtain a $0+1$-dimensional super quantum mechanics (SQM) with eight supercharges and gauge group ${\rm U}(n)$. A bound state with charge $nq_{0}$ exists if it is possible to Higgs the ${\rm U}(n)$ gauge group to a single ${\rm U}(1)$. For this to be possible, the SQM obtained from the wrapped D3-branes must have sufficiently many scalar fields. For a special Lagrangian cycle $\Gamma_{0}$ the number of scalar deformation modes is counted by $b_{1}(\Gamma_{0})$. The SQM has enough scalar modes to realise the Higgsing ${\rm U}(n)\to{\rm U}(1)$ if $b_{1}(\Gamma_{0})\geq 3$. [^20] To see this, we notice that the case $b_{1}(\Gamma_{0})=3$ corresponds to $\Gamma_{0}$ being topologically a $T^{3}$ in which case all sixteen supercharges of the parent 4d $\mathcal{N}=4$ ${\rm U}(n)$ SYM theory living on the worldvolume theory of the $n$ D3-branes are preserved. In this case the fields of the resulting SQM form complete multiplets of the super-algebra with 16 supercharges. As is well-known from the case of D0-branes, the scalar fields in the multiplets of $\mathcal{N}=16$ are sufficient to completely break ${\rm U}(n)\to{\rm U}(1)$. This requirement translates into a condition on $C_{0}$ since due to the fibration structure of $\Gamma_{0}$, $b_{1}(\Gamma_{0})$ is bounded from below by the first Betti number of the base, $b_{1}(\Gamma_{0})\geq b_{1}(C_{0})$. For the curve $C_{0}$, the first Betti number is given by its genus $g(C_{0})$, $b_{1}(C_{0})=2g(C_{0})$. Hence, for the special Lagrangian cycles $\Gamma_{0}$, bound states with charge $nq_{0}$ exist if the curve $C_{0}\in H_{2}(Z)$ has genus $g(C_{0})\geq 1$. [^21] Here, we include the case $g(C_{0})=1$ since in this case the $S^{1}$-fibration over $C_{0}$ is trivial such that $b_{1}(\Gamma_{0})=b_{1}(C_{0})+1=3$.
@@ 3.2 Tensionless Strings
where $\partial=(\partial/\partial z){\rm d}z$. Simple BPS solutions to these equations of motions are obtained for holomorphic profiles ${\bar{\partial}}t=0$. The Einstein equations then fix the warp factor to be
@@
[…]
@@
$\frac{T_{\mathbf{e}_{k_{1}}}}{m_{\rm tower}^{2}}\to\text{const.}\,$, (3.30)